$SPAD/src/input knot.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

\getchunk{license}
)cl all
-- Create a (p,q) torus-knot with radius around the curve of r.
torusKnot(p,q,r) ==
a := 1
b := 1
ab := a+b
fac := 2*ab/(1.1*ab-sin(q*t))
Title := concat("Torus Knot of Type (", concat(p::String,
                concat(",", concat(q::String, ")"))))
Curve := curve(a*fac*cos(p*t), a*fac*sin(p*t), b*fac*cos(q*t))
r = 0 => draw(Curve, t=0..2*%pi, title == Title)
draw(Curve, t=0..2*%pi, tubeRadius == r, title == Title)
torusKnot(2, 3, 0.3)

———-

References^{[1] nothing}

4